In a winch assembly plant at Curt Manufacturing, the winch assembly work cell for the winch mount handle has a demand of 200 units an hour. Two feeder cells supply the two winch component parts CRT31005 and CRT 31006 to the assembly cell (one CRT31005 and one CRT31006 for each winch mount handle). Standard containers will hold 20 CRT31005’s or 10 CRT31006’s. It takes 10 minutes to fill up a container with CRT31005’s and 20 minutes to full up a container with CRT31006’s. The transit time to the assembly work cell is 5 minutes for both CRT31005 and CRT31006; the policy variable set by management that reflects possible inefficiency in the system is 0. . Set up a Kanban control system for the assembly process.
Answer #1

Where K = Number of Kanbans required

d = average production rate per day

w = waiting time in fractions of a day

p = processing time in fractions of a day

c = capacity of container

= efficiency of process

For CRT31005:

d = 200*24 = 4800

w = 10/(60*24) = 0.00694

p = 5/(60*24) = 0.0347

c = 20

= 0

K = = 10 Kanbans

For CRT31006:

d = 200*24 = 4800

w = 20/(60*24) = 0.01388

p = 5/(60*24) = 0.0347

c = 10

= 0

K = = 23.3 Kanbans (rounding up to 24)

Thus,

CRT31005 requires 10 Kanbans

CRT31006 requires 24 Kanbans